OpenStudy (anonymous):

Enter in standard form the equation of the line passing through the given point and having the given slope. B(2, - 5), m =

OpenStudy (anonymous):

m=-2/3

OpenStudy (ajprincess):

\(y-y_1)=m(x-x_1)\) \(y-(-5)=-2/3(x-2)\) \(3(y+5)=-2(x-2)\) standard form is ax+by=c Can u do it?

OpenStudy (anonymous):

3y+15=-2x+4

OpenStudy (ajprincess):

3y+15=-2x+4 1.Add 2x to both sides 2.subtract 15 both sides.

OpenStudy (anonymous):

3y+2x=-11

OpenStudy (ajprincess):

ya perfect:)

OpenStudy (anonymous):

thanks so much

OpenStudy (ajprincess):

yw:)